A particle moves along the x-axis so that at time t≥0 its position is given by x(t) = -t^3 + 14t^2 25t. Determine the velocity of the particle at t = 2.


1
Expert's answer
2021-02-10T01:44:48-0500

"Given:\;\;Particle's\;position\;is\;given\;by \to\;\;x(t)=-t^3+14t^2+25t"

"To\;find:\;Velocity\;of\;the\;particle\;at\;t=2,\;i.e.,\;x'(2)"


"We\;know\;that\;-\;\;v(t)=\dfrac{d}{dt}(x(t))=\dfrac{d}{dt}(-t^3+14t^2+25t)"


"=-\dfrac{d}{dt}(t^3)+14\dfrac{d}{dt}(t^2)+25\dfrac{d}{dt}(t)"


"=-(3t^2)+14(2t)+25(1)"

"=-3t^2+28t+25"


"\therefore x'(t)=-3t^2+28t+25\;and\;\;x'(2)=-3(2^2)+28(2)+25=-12+81=\mathbf{69}"